OpenStudy (anonymous):

f(x)=3x+9, g(x)=5x^2 find (f+g)(x) help please

OpenStudy (aum):

(f+g)(x) = f(x) + g(x) = 3x + 9 + 5x^2 = 5x^2 + 3x + 9